OpenStudy (anonymous):

Move everything away from y by using the additive inverse/opposite 2x + y = 11

OpenStudy (muhammad_nauman_umair):

2x-2x+y=11-2x y=11-2x

OpenStudy (anonymous):

i mean y=-2x+11